Calculate the mole fraction of H2SO4 in 8%(w/w) aqueous H2SO4 ( molar mass: 98 g/mol, 18g/mol respectively


1
Expert's answer
2021-03-04T06:32:53-0500

8%(w/w) aqueous H2SO4 solution means that 100 g of solution contains 8 g of the acid

Mass of water: 100 – 8 = 92 (g)

Moles of each:

H2SO4: 8 g / 98 g/mol = 0.082 mol

H2O: 92 g / 18 g/mol = 5.111 mol

Total moles: 5.193 mol

Mole fraction:  

H2SO4: 0.082 / 5.193 = 0.0157 or 1.57%

H2O: 100 – 1.57 = 98.43%